Find the interval of convergence of P∞
n=0
x
n
n+4 .
1
Expert's answer
2020-06-16T19:17:31-0400

Given series is "\sum^\infin_{n=0} \frac{x^n}{n+4}."

Let "a_n = \frac{1}{n+4}"

"\implies a_{n+1} = \frac{1}{n+5}"

So, Radius of convergence is "R= \lim_{n \to \infin} \frac{a_n}{a_{n+1}} = \lim_{n \to \infin} \frac{n+5}{n+4} = \lim_{n \to \infin} \frac{1+5/n}{1+4/n} =1" .

So, series is convergence on "(-1,1)" is convergent and on "(-\infin,-1) \cup(1,\infin)" series is divergent.

Now, at "x=1" series becomes "\sum^\infin_{n=0} \frac{1}{n+4}" which is divergent by p-test.

At "x=-1" series becomes "\sum^\infin_{n=0} \frac{(-1)^n}{n+4}", which is convergent by Leibniz's test because "<\frac{1}{n+4}>" is decreasing sequence and"\frac{1}{n+4} \to 0" as "n\to \infin".

So interval of convergence is "[-1,1)" .